OpenStudy (ttp):

CHECK MY ANSWER!!! I THINK IT'S OPTION A Tiffany kicks a soccer ball off the ground and in the air, with an initial velocity of 28 feet per second. Using the formula H(t) = -16t2 + vt + s, what is the maximum height the soccer ball reaches? 15.6 feet 14.1 feet 13.5 feet 12.3 feet

OpenStudy (dqswag):

that "s" represents the INITIAL HEIGHT of the soccer ball. Obviously it's not going to be a very large number, since most of us could not kick a soccer ball that was more than about 5 feet in the air, right? If v = initial velocity = 28 (ft/sec), then your equation becomes APPROXIMATELY H(t) = -16*t^2+28t+5. Note that you didn't include the t^2 earlier, so I've fixed that. Now, if you'd graph H(t) = -16*t^2+28t+5, you could see pretty easily what the maximum height will be. Sound OK?

OpenStudy (dqswag):

15.6